OpenStudy (anonymous):

Complete the equations of the system in slope-intercept form. Use a decimal for the slope if necessary. Line 1 x y 0 3 1 5 Line 2 x y –1 1 –2 –1 Line 1: y = x + Line 2: y = x +

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

First, find the slope. \(\sf m = \dfrac{y_2-y_1}{x_2-x_1}\) Line 1: (0, 3), (1, 5) x1 y1 x2 y2 So we have: \(\sf m = \dfrac{5-3}{1-0}\) Simplify

OpenStudy (anonymous):

2/1

OpenStudy (igreen):

Yep, or 2. Now plug it into point-slope form along with any of the two points. Let's use (1, 5). Put it in point-slope form first. \(\sf y - y_1 = m(x - x_1)\) Where \(\sf y_1\) is the y-value of the point, \(\sf x_1\) is the x-value of the point, and \(\sf m\) is the slope. So we have: \(\sf y - 5 = 2(x - 1)\) Distribute 2 into the parenthesis, can you do that?

OpenStudy (anonymous):

y-5=2x-2

OpenStudy (igreen):

Yep, now add 5 to both sides.

OpenStudy (anonymous):

y=2x-3

OpenStudy (igreen):

Not quite, -2 + 5 = ?

OpenStudy (anonymous):

-7

OpenStudy (igreen):

Check again.

OpenStudy (anonymous):

but your adding negatives

OpenStudy (anonymous):

+3

OpenStudy (anonymous):

@igreen